28. Juni 2007. U. Schoenwaelder; http://www.math.rwth-aachen.de/∼Ulrich.Schoenwaelder HB = Hochschulbibl. RWTH, HBZ = http://www.hbz-nrw.de/ (HBZ-CD-ROM Online), MB = Mathematikbibl., DB = Didaktikbibl. (Winter), EB = Erziehungswiss. Bibl., FH = Bibl. Fachhochschule Aachen, FL = Fernleihe, IB = Informatikbibliothek, IB Nr. = Institutsbibliothek Nr., LB = HB– Lehrbuchsammlung, LS = HB–Lesesaal

Literatur zur Galoistheorie

[1] Jörg Bewersdorf. Galois Theory for Beginners: a Historical Perspective. Student Mathematical Library 35. AMS, 2006.

[2] Jörg Bewersdorff. Algebra für Einsteiger. Von der Gleichungsauflösung zur Galois-Theorie. Vieweg, 2002. ISBN 3-528-03192-1. HBZ.

[3] H. M. Edwards. Essays in Constructive Mathematics. Springer-Verlag, 2005. ISBN 0-387-21978-1. HBZ. Uses finite algorithms. The topics covered derive from classic works of nineteenth century mathematics: Galois’ theory of algebraic equations, Gauss’ theory of binary quadratic forms, Abel’s theorem about integrals of rational differentials on algebraic curves (Newton’s polygon), fundamental theorem of algebra, factorization of polynomials over an algebraic number field, spectral theorem for symmetric matrices.

[4] J.-H. Eschenburg and L. Hefendehl-Hebeker. Die Gleichung 5. Grades: Ist Mathematik erzählbar? Mathematische Semesterberichte, 47(2):193–220, 2000. HB: Z1538 (ZNT). Ikosaedergruppe.

[5] M. H. Fenrick. Introduction to the Galois Correspondence. Birkhäuser, 1991. MB: 16428.

[6] L. Gaal. Classical Galois Theory with Examples. Chelsea Publ. Co., 1971, ^{2}1973.

[7] D. J. H. Garling. A Course in Galois Theory. Cambridge Univ. Press, 1986. MB: 13523.

[8] John M. Howie. Fields and Galois Theory. Springer undergraduate mathematics series. Springer, 2006. ISBN 978-1-85233-986-9; 1-85233-986-1.

[9] R. Bruce King. Beyond the Quartic Equation. Birkhäuser Boston, 1996. ISBN 0-8176-3776-1, 3-7643-3776-1.MB: 17888.

[10] M. Kuga. Galois Dream: Group Theory and Differential Equations. Birkhäuser, 1993, ^{3}1997. MB: A 16898.

[11] E. Kunz. Algebra. Vieweg–Studium: Aufbaukurs Mathematik 43. Vieweg, Braunschweig, 1991. MB: 16021, HB: Bf 9311, Bf 9311+2 (1994). Wohlmotivierte Einführung in Körper– und Galois–Theorie.

[12] Theophil Lambacher. Eine Zahlenfolge und antike geometrische Probleme. Praxis der Mathematik, 10:1–4, 1968. MB: Z 101. Winkeldreiteilung und dritte Wurzel (Delisches Problem) per rationaler Approximation; geometrische Konstruktion.

[13] M. Neubrand. Didaktik – Zahlen – Algebra: mathematisch-didaktische Überlegungen am Fundamentalsatz der Algebra. Texte zur mathematisch-naturwissenschaftlich-technischen Forschung und Lehre 17. Franzbecker, 1985. ISBN 3-88120-014-2. FL: UB Bonn U 85/6748. Habilitationsschrift Bonn, 1984. Vgl. [14].

[14] M. Neubrand. Hochschuldidaktische Überlegungen zum Fundamentalsatz der Algebra. J. für Mathematikdidaktik, 6(1):45– 66, 1985. HB: Z5899. FSA via Galoistheorie als Leitidee für eine Algebra–Vorlesung. Vgl. [13].

[15] Barbara L. Osofsky. Nice polynomials for introductory Galois theory. Mathematics Magazine, 72(3):218–222, 1999. MB: Z 167.

[16] L. Toti Rigatelli. Evariste Galois (1811–1832). Birkhäuser, 1996. MB: 18065.

[17] Regina Röhner. Der Provokateur. Historischer Roman. Bestellnr. 639 446 3. Mitteldeutscher Verlag, 1989. Galois 1811– 1832. Hinweis in Math. Schule (HB: Z5724) 29:1 (1991), 77.

[18] J. Rotman. Galois Theory. Universitext. Springer–Verlag, 1990. MB: 15404, HB: Bb 1784.

[19] Leila Schneps, editor. Galois Groups and Fundamental Groups, Mathematical Sciences Research Institute Publications 41. Cambridge Univ. Press. ISBN 0-521-80831-6. Eight expository articles focus on recent developments.

[20] Leila Schneps, editor. The Invers Galois Problem, Moduli Spaces and Mapping Class Groups (Conf. ”Geometry and Arithmetic of Moduli Spaces“, Luminy), Geometric Galois actions 2. Cambridge Univ. Press, 1997. ISBN 0-521-59641-6.

[21] J.-P. Serre. Topics in Galois Theory. Research Notes in Mathematics 1. Jones and Bartlett Publ., 1992. Fortgeschritten: Umkehrproblem der Galoistheorie.

[22] I. Stewart. Galois Theory. Chapman and Hall Mathematical Series. Chapman and Hall, 1973, ^{3}2004. MB: 7107. Sehr gut lesbar. 3rd edition: reorganization of the material places the concrete before the abstract, but the substance of the book remains the same.

[23] G. Stroth. Algebra. Einführung in die Galoistheorie. de Gruyter Lehrbuch. de Gruyter, 1998. Beispiele.

[24] Jean-Pierre Tignol. Galois’ Theory of Algebraic Equations. Longman, 1988. HB: Bf9050. ISBN 0-582-00290-7, 0-470-20919-4.

[25] Helmut Völklein, editor. Progress in Galois Theory: Proceedings of John Thompson’s 70th Birthday Conference (2002, Gainesville, Fla), Developments in mathematics 12. New York: Springer, 2005. 0-387-23533-7; 0-387-23534-5 e-book. HBZ.